# Learning Efficient Representations of Neutrino Telescope Events

### v1

@@ 1 Introduction
Neutrino telescopes search for rare interactions caused by neutrinos, an elusive particle central to many open questions in the Standard Model of particle physics, such as the origin of their masses and the matter anti-matter asymmetry of the universe. The leading detector in operation today, the IceCube Neutrino Observatory, comprises 5,160 optical modules (OMs) arranged on strings deep within the clear ice of the Antarctic glacier (see Aartsen et al. (2017)). Each OM is designed to precisely record the arrival times of photons, forming a photon arrival time distribution (PATD). The data gathered by IceCube are archived as “events,” which are triggered when a specific number of coincident light detections occur between OMs in neighboring strings, signaling the presence of a charged particle. Figure 1 is an artistic representation of the photons produced in a neutrino interaction and their subsequent detection in the OMs.
@@
A neutrino telescope typically classifies events into categories based on the morphology of the light detected: cascade-like events, identifiable by their approximately spherical spatial distribution; and track-like events, characterized by their elongated, linear spatial morphology. In IceCube, events are recorded at an approximate rate of 3000 per second, constituting a large data rate. Moreover, comparable or even larger optical neutrino telescopes are either under construction or planned, such as KM3NeT (Aiello et al., 2019), Baikal-GVD (Avrorin et al., 2018), P-ONE (Agostini et al., 2020), IceCube-Gen2 (Aartsen et al., 2021), TRIDENT (Ye et al., 2022), and HUNT (Huang et al., 2023), signalling a need to develop methods for efficiently handling neutrino telescope data.
The accurate reconstruction and prediction of the interacting particle’s physical properties—including its energy, direction, and type— from the PATDs recorded by each OM, is essential for informing and facilitating downstream physics analyses. ML techniques have successfully been applied to this task in recent years, notably in neutrino telescopes like IceCube (Choma et al., 2018; Abbasi et al., 2021; Huennefeld et al., 2021; Abbasi et al., 2022; Yu et al., 2023; Eller, 2023; Jin et al., 2024; Zhu et al., 2024) and KM3NeT (De Sio, 2019; Aiello et al., 2020; Reck et al., 2021; Guderian, 2022; Mauro & de Wasseige, 2023). Nevertheless, a significant challenge in ML-based reconstruction techniques remains: the incorporation of the full-timing information provided by the PATDs of the OMs, which are characterized by high dimensionality and variable lengths. Interaction events may last several $\mu s$, yet ns-scale timing resolution is needed on the PATDs for most downstream physics analyses. Furthermore, as illustrated by OM#1 in Figure 1, some OMs near the interaction point can detect up to tens of thousands of photons in the highest-energy and brightest events. On the other hand, many further OMs record only a few photons, resulting in highly irregular events; illustrated in the OM#2 distribution on the same figure.
To address the challenges of working with this data, previous studies have employed summary statistics (Abbasi et al., 2021), using key variables that summarize the PATD of a given OM. This approach yields greater efficiency and flexibility, as each OM has a fixed-size description of the timing distribution, allowing events to be formatted as 2D images with multiple features per pixel. However, the summary statistics lose a substantial amount of information present in the original PATD. An alternative method involves parameterizing the PATD by fitting an asymmetric Gaussian mixture model (AGMM). In this study, we implemented a basic AGMM to serve as a comparison, drawing inspiration from Huennefeld et al. (2021). While this retains significantly more information than summary statistics, the optimization process can be slow and prone to failure, with a strong dependence on hyperparameters, e.g., the number of Gaussian components. In this paper, we take a deep learning approach for the first time by introducing a transformer-based variational autoencoder (VAE), termed om2vec, which converts PATDs into a fixed-size compact latent representation. om2vec is trained to reconstruct the input PATD as readout by OMs, using only the information from its learned latent space. Our approach provides a number of significant benefits over previous methods:
  • om2vec generally retains more information compared to the AGMM approach, giving it greater flexibility in representing more complex PATDs.
  • om2vec demonstrates greater reliability by avoiding optimization failures.
  • om2vec is orders of magnitude faster than the AGMM, especially when scaling to larger latent dimensions. It can also benefit more from GPU acceleration and parallel processing.
  • om2vec is less dependent on hyperparameters.
We argue that these advantages make this the first “one-size-fits-all” neutrino event representation learner, making it suitable for deployment at the earliest stages of experimental data collection. With a robust latent representation of each OM’s PATD, neutrino events can easily be arranged into image-like data. This will serve to circumvent the challenges inherent in raw neutrino telescope data, and facilitate the rapid adoption of more sophisticated ML techniques from the image-processing community into the neutrino physics field. Source code, datasets, and pre-trained checkpoints can be found on GitHub.
@@ 2 Architecture
Since their inception, VAEs have been extensively utilized for learning effective data representations and generative modeling (Kingma & Welling, 2022). More recently, transformers have demonstrated exceptional performance in handling sequential data (Vaswani et al., 2017). Given the time series nature of our dataset, we incorporate transformer layers to enhance feature learning capabilities. Our model employs the standard encoder-decoder structure, as illustrated in Figure 2. We assume a Gaussian prior for the latent space and train the encoder to learn the parameters $\mu$ and $\sigma$, with some added random noise $\epsilon$. The re-parameterization trick is utilized to construct the latent representation $z$, a vector of user-defined length referred to as the latent dimension. This technique guarantees that the latent space remains continuous, and that similar representations within this space reconstruct to similar PATDs.
@@
The input data PATDs are formatted as one-dimensional time series sequences of equal length, with each element corresponding to a specific time in sequence. Each element features a single attribute: the number of photon hits occurring during that particular time window. We first expand the feature space with an input embedding layer. The encoder and decoder are constructed by stacking blocks. Importantly, we employ a hybrid approach that operates across both the feature and sequence dimensions. Transformer layers process the feature dimension, followed by a feed-forward network that down-samples or up-samples the sequence length in the encoder and decoder, respectively. Linear layers are then used to flatten the feature dimension into latent representations, as well as to generate output logits after the decoder. We opted for a learnable memory embedding for the transformer decoder layers, ensuring that the decoder portion of the architecture remains entirely independent of the encoder. This design allows for future utilization directly on latent representations without requiring outputs from the encoder or the input PATDs themselves.
3 Dataset and Training Details
@@
Simulations play a key role in neutrino physics experiments, facilitating preliminary testing that informs experiment development and subsequent physics analyses. As such, the field has developed sophisticated simulation tools capable of generating events for incorporation into our training datasets. In this paper, we present results trained and tested on simulated events, but emphasize that the methodology can be readily applied to real detector data. We employ the open-source simulation toolkit Prometheus (Lazar et al., 2024) to emulate an IceCube-like array of OMs within Antarctic ice. We generate four datasets of events based on the different neutrino interaction types, corresponding to the distinct flavors of the neutrino, $\nu$. Specifically, these interactions are the charged-current electron neutrino ($\nu_{e}$ CC), muon neutrino ($\nu_{\mu}$ CC), and tau neutrino ($\nu_{\tau}$ CC) interactions, alongside the neutral-current ($\nu$ NC) interactions. Notably, the $\nu_{e}$ CC, $\nu_{\tau}$ CC, and $\nu$ NC interactions produce predominantly spherical cascade-like events, while the $\nu_{\mu}$ CC interactions result in long track-like events. Additionally, as further detailed in Section 4, the PATDs generated by $\nu_{\tau}$ CC interactions are of particular interest to physicists due to their potential to exhibit a smoking-gun “double-bang” signature of astrophysical neutrinos.
@@
Table 1: Model and training hyperparameters.
PARAMETER VALUE
Input Sequence Length 6400
Latent Dimension [32, 64, 128]
Embedding Feature Size 32
No. Trainable Parameters 15M
Batch Size 1024
As previously mentioned, each PATD in the datasets is pre-processed into a fixed-length vector by binning photon hits over time. A total of 6400 bins, each 1 ns wide, represent the first 6.4 $\mu s$ of hits—a readout time frame consistent with IceCube’s OMs (Aartsen et al., 2017). The choice of binning and time window can be adapted to suit the specifications of other neutrino telescopes. Any remaining hits are accumulated in a final overflow bin. The hit counts in each bin are then log-normalized, as they can be extremely large for high-energy neutrino interactions. A table of model and training hyperparameters can be found in Table 1.
The training loss combines reconstruction loss and a KL divergence regularization loss on the latent space. The reconstruction loss is computed as the Poisson negative log-likelihood based on the input PATD. Thus, the network is trained to predict the underlying probability density function that generates the input PATD. The KL divergence loss term is scaled by a $\beta$ factor that follows a cyclic cosine function during training, peaking at a hyperparameter value set at $10^{-5}$.
@@ 4 Results
where $\mu$ and $\sigma$ are the mean and standard deviation and $r$ is the asymmetry parameter, as introduced in (Huennefeld et al., 2021). Each asymmetric Gaussian is assigned a weighting factor, $w$, and then summed. Consequently, each component of the mixture model introduces four parameters: $\mu$, $\sigma$, $r$, and $w$. The mixture model is initially fit using $k$-means clustering to obtain the starting parameters, followed by solving a bounded-constrained optimization problem that minimizes the negative log-likelihood. However, the fit is only performed when the number of photon hits (the length of $x$) is greater than the number of Gaussian components. Otherwise, the true photon hit times are stored.
@@ 4.1 Timing Distribution Reproducibility
To assess the reconstruction capability, we employ the Jensen-Shannon (JS) distance (Lin, 1991; Nielsen, 2019), which is derived from the KL divergence and is defined as follows:
@@
Figure 5: An example PATD of a “double-bang” signature, characterized by two distinct peaks. om2vec is able to reconstruct both peaks, while the AGMM is noticeably less accurate.
In order to give a more intuitive sense of the performance of om2vec, we examine individual PATDs and their reconstructed PDFs. This is especially interesting for exotic cases such as the “double-bang” signature produced by $\nu_{\tau}$ interactions, as mentioned earlier. Figure 5 illustrates an example of this type of PATD, distinguished by its two distinct peaks. It is evident that om2vec reconstructs the bimodal structure of the timing distribution with greater precision than the AGMM method, despite the overall JS distance difference between the two methods being minor. As this is an extremely large PATD ($\sim$100,000 photons), this is likely due to the dominant statistics contained in the first peak. However, from a physics perspective, retaining information about both peaks is crucial for downstream tasks that attempt to isolate $\nu_{\tau}$ interactions, a smoking-gun indication of astrophysical neutrinos.
@@
  • SSCNN (Full): a sparse submanifold convolutional neural network (SSCNN) that utilizes full timing information in a 4D CNN, as described in Yu et al. (2023).
@@
In Figure 6, we show the neutrino angular resolution (the angular difference between the true and predicted neutrino direction), as a function of the true neutrino energy. We find that the SSCNN models using the full timing information and the latent representations perform similarly, though as expected, the full 4D model is able to achieve a slightly better resolution. Notably, SSCNN sacrifices some performance to remain computationally feasible on higher-dimensional sparse data (Graham & van der Maaten, 2017). Thus, we observe that the dense CNN approach with latent representations performs just as well or slightly better than the SSCNN with full-timing information, depending on the true neutrino energy. This CNN approach is made possible by the fixed-length latent representations. Importantly, we note that this new flexibility allows for the easy adaptation of more powerful models, such as vision transformers (Dosovitskiy et al., 2021), for angular reconstruction.
@@ 5 Conclusions and Future Directions
Availability: Source code, datasets, and pre-trained checkpoints for om2vec on Prometheus events is made available on GitHub. Additionally, as an immediate next step, om2vec is being integrated into GraphNeT (Søgaard et al., 2023), an open-source deep learning pipeline widely utilized by various collaborations in the neutrino telescope community.

### v2

@@ 1 Introduction
Neutrino telescopes search for rare interactions caused by neutrinos, an elusive particle central to many open questions in the Standard Model of particle physics, such as the origin of their masses and the matter anti-matter asymmetry of the universe. The leading detector in operation today, the IceCube Neutrino Observatory, comprises 5,160 optical modules (OMs) arranged on strings deep within the clear ice of the Antarctic glacier (see [1]). Each OM is designed to precisely record the arrival times of photons, forming a photon arrival time distribution (PATD). The data gathered by IceCube are archived as "events," which are triggered when a specific number of coincident light detections occur between OMs in neighboring strings, signaling the presence of a charged particle. Figure 1 is an artistic representation of the photons produced in a neutrino interaction and their subsequent detection in the OMs. A neutrino telescope typically classifies events into categories based on the morphology of the light detected: cascade-like events, identifiable by their approximately spherical spatial distribution; and track-like events, characterized by their elongated, linear spatial morphology. In IceCube, events are recorded at an approximate rate of 3000 per second, constituting a large data rate. Moreover, comparable or even larger optical neutrino telescopes are either under construction or planned, such as KM3NeT [2], Baikal-GVD [3], P-ONE [4], IceCube-Gen2 [5], TRIDENT [6], and HUNT [7], signaling a need to develop methods for efficiently handling neutrino telescope data.
The accurate reconstruction and prediction of the interacting particle’s physical properties—including its energy, direction, and type— from the PATDs recorded by each OM, is essential for informing and facilitating downstream physics analyses. A significant challenge in this task is the incorporation of the full-timing information provided by the PATDs of the OMs, which are characterized by high dimensionality and variable lengths. Interaction events may last several $\mu s$, yet ns-scale timing resolution is needed on the PATDs for most downstream physics analyses. Furthermore, as illustrated by OM#1 in Figure 1, some OMs near the interaction point can detect up to tens of thousands of photons in the highest-energy and brightest events. On the other hand, many further OMs record only a few photons, resulting in highly irregular events; illustrated in the OM#2 distribution on the same figure. In this paper, we introduce a transformer-based variational autoencoder (VAE), termed om2vec, which encodes PATDs into a fixed-size compact latent representation. om2vec is trained to reconstruct the input PATD as readout by OMs, using only the information from its encoded latent space. In this sense, it learns an efficient and more flexible representation of the data recorded in each OM. Our approach provides several significant advantages over existing methods, including improved information retention, greater reliability, faster processing speed, and reduced hyperparameter dependence. We also demonstrate that utilizing transformers for this particular problem is particularly effective, by comparing it to a baseline fully-connected network. We argue that these advantages make this the first “one-size-fits-all" neutrino event representation learner, making it suitable for deployment at the earliest stages of experimental data collection. This will serve to circumvent the challenges inherent in raw neutrino telescope data, and facilitate the rapid adoption of more sophisticated ML techniques from the image-processing community into the neutrino physics field. Source code, datasets, and pre-trained checkpoints can be found on GitHub.
@@
1.1 Related Works
Machine learning techniques have successfully been applied to neutrino reconstruction tasks in recent years, notably in neutrino telescopes such as IceCube [8, 9, 10, 11, 12, 13, 14, 15] and KM3NeT [16, 17, 18, 19, 20]. Previous studies have employed various approaches to address the challenges of working with neutrino telescope data. One approach uses summary statistics [9], utilizing key variables that summarize the PATD of a given OM. This method yields greater efficiency and flexibility, as each OM has a fixed-size description of the timing distribution, allowing events to be formatted as 2D images with multiple features per pixel. However, the summary statistics lose a substantial amount of information present in the original PATD. An alternative method involves parameterizing the PATD by fitting an asymmetric Gaussian mixture model (AGMM). In this study, we implemented a basic AGMM to serve as a comparison, drawing inspiration from [10]. While this retains significantly more information than summary statistics, the optimization process can be slow and prone to failure, with a strong dependence on hyperparameters, e.g., the number of Gaussian components. Our transformer-based VAE approach builds upon these previous methods while addressing their limitations.
@@ 2 Architecture
Since their inception, VAEs have been extensively utilized for learning effective data representations and generative modeling [21]. More recently, transformers have demonstrated exceptional performance in handling sequential data [22]. Given the time series nature of our dataset, we incorporate transformer layers to enhance feature learning capabilities. Our model employs the standard encoder-decoder structure, as illustrated in Figure 2. We assume a Gaussian prior for the latent space and train the encoder to learn the parameters $\mu$ and $\sigma$, with some added random noise $\epsilon$. The re-parameterization trick is then utilized to construct the latent representation $z$ while maintaining proper gradient flow to these parameters. A key property of VAEs over regular autoencoders is their continuous latent space, meaning that similar representations within this space correspond to similar reconstructed PATDs.
@@
The input data PATDs are formatted as one-dimensional time series sequences of equal length, with each element corresponding to a specific time in sequence. Each element features a single attribute: the number of photon hits occurring during that particular time window. We first expand the feature space with an input embedding layer. The encoder and decoder are constructed by stacking blocks. Importantly, we employ a hybrid approach that operates across both the feature and sequence dimensions. Transformer layers process the feature dimension, followed by a feed-forward network that down-samples or up-samples the sequence length in the encoder and decoder, respectively. Linear layers are then used to flatten the feature dimension into latent representations, as well as to generate output logits after the decoder. We use a simple vector of learnable parameters, the "memory embedding," which acts as the memory input for the transformer decoder layers. After the final linear layer, the outputs are fed through the softmax function to obtain a properly normalized probability density. We interpret the final softmax activation’s output as the probability of detecting a photon hit in each timing bin.
3 Datasets
@@
Simulations play a key role in neutrino physics experiments, facilitating preliminary testing that informs experiment development and subsequent physics analyses. As such, the field has developed sophisticated simulation tools capable of generating events for incorporation into our training datasets. In this paper, we present results trained and tested on simulated events, but emphasize that the methodology can be readily applied to real detector data. We employ the open-source simulation toolkit Prometheus [23] to emulate an IceCube-like array of OMs within Antarctic ice. We generate four datasets of events based on the different neutrino interaction types, corresponding to the distinct flavors of the neutrino, $\nu$. Specifically, these interactions are the charged-current electron neutrino ($\nu_{e}$ CC), muon neutrino ($\nu_{\mu}$ CC), and tau neutrino ($\nu_{\tau}$ CC) interactions, alongside the neutral-current ($\nu$ NC) interactions. Notably, the $\nu_{e}$ CC, $\nu_{\tau}$ CC, and $\nu$ NC interactions produce predominantly spherical cascade-like events, while the $\nu_{\mu}$ CC interactions result in long track-like events. Additionally, as further detailed in Section 5, the PATDs generated by $\nu_{\tau}$ CC interactions are of particular interest to physicists due to their potential to exhibit a smoking-gun “double-bang" signature of astrophysical neutrinos.
@@
As previously mentioned, each PATD in the datasets is pre-processed into a fixed-length vector by binning photon hits over time. A total of 6400 bins, each 1 ns wide, represent the first 6.4 $\mu s$ of hits—a readout time frame consistent with IceCube’s OMs [1]. The choice of binning and time window can be adapted to suit the specifications of other neutrino telescopes. Any remaining hits are accumulated in a final overflow bin. The hit counts in each bin are then log-normalized, as they can be extremely large for high-energy neutrino interactions.
4 Proposed Methods
In this section, we provide a more detailed definition of the objective function for this problem. As previously discussed, the conceptual goal of om2vec is to learn how to encode a PATD into a representation vector and subsequently decode this representation back into its corresponding probability density function. Mathematically, the reconstruction loss is defined as
$\mathcal{L}_{reco}(\theta)=-\log\mathcal{P}(\mathbf{n}|\theta)=\sum_{i=1}^{N_{\text{bins}}}\left[-n_{i}\log(\lambda_{i}(\theta))+\lambda_{i}(\theta)\right]$ (4.1)
where:
[…]
Given the predicted PDF from the network (after softmax), and the normalized true PDF from the input PATD, this loss function takes the negative logarithm of the summed Poisson likelihoods, representing the probability of registering the observed number of hits in each bin, based on the predicted hit counts.
The total training loss combines reconstruction loss and a KL divergence regularization loss on the latent space. This regularization term is needed to ensure that the learned latent space aligns with the prior normal distribution. It is also scaled by a $\beta$ factor that follows a cyclic cosine function during training, peaking at a hyperparameter value set at $10^{-5}$. A batch size of 1024 is used during training.
@@ 5 Results
where $\mu$ and $\sigma$ are the mean and standard deviation and $r$ is the asymmetry parameter, as introduced in [10]. Each asymmetric Gaussian is assigned a weighting factor, $w$, and then summed. Consequently, each component of the mixture model introduces four parameters: $\mu$, $\sigma$, $r$, and $w$. The mixture model is initially fit using $k$-means clustering to obtain the starting parameters, followed by solving a bounded-constrained optimization problem that minimizes the negative log-likelihood. However, the fit is only performed when the number of photon hits (the length of $x$) is greater than the number of Gaussian components. Otherwise, the true photon hit times are stored.
@@ 5.1 Timing Distribution Reproducibility
To assess the reconstruction capability, we employ the Jensen-Shannon (JS) distance [24, 25], which is derived from the KL divergence and is defined as follows:
@@
Figure 5: An example PATD of a “double-bang" signature, characterized by two distinct peaks. om2vec is able to reconstruct both peaks, while the AGMM is noticeably less accurate.
In order to give a more intuitive sense of the performance of om2vec, we examine individual PATDs and their reconstructed probability density functions. This is especially interesting for exotic cases such as the “double-bang” signature produced by $\nu_{\tau}$ interactions, as mentioned earlier. Figure 5 illustrates an example of this type of PATD, distinguished by its two distinct peaks. It is evident that om2vec reconstructs the bimodal structure of the timing distribution with greater precision than the AGMM method, despite the overall JS distance difference between the two methods being minor. As this is an extremely large PATD ($\sim$100,000 photons), this is likely due to the dominant statistics contained in the first peak. However, from a physics perspective, retaining information about both peaks is crucial for downstream tasks that attempt to isolate $\nu_{\tau}$ interactions, a smoking-gun indication of astrophysical neutrinos.
Table 1: Average JS distance and forward-pass FLOPs for the different om2vec models tested.
MODEL JSD FLOPs
Fully-connected 0.3545 $2.899\times 10^{7}$
Transformers (1 block) 0.2340 $1.299\times 10^{9}$
Transformers (3 blocks) 0.2177 $\bf{1.882\times 10^{9}}$
We further evaluated three different model architectures, all employing 64 latent parameters, to assess their impact on the average JS distance. As a baseline case, we trained a model using only fully-connected layers. The other two architectures incorporated transformers, as it is described in Section 2, with varying numbers of encoder and decoder blocks. A summary of the results is provided in Table 1. We note that the default architecture used in previous experiments features three encoder-decoder blocks and is highlighted in bold. Comparing the fully connected model to the transformer-based architectures, we observe a substantial improvement in reconstruction performance when incorporating transformer layers. However, this enhancement comes at a considerable computational cost, with the number of FLOPs required for a forward pass increasing by approximately two orders of magnitude. Furthermore, adding additional transformer encoder-decoder layers yields a slight further improvement in reconstruction performance.
@@
  • SSCNN (Full): a sparse submanifold convolutional neural network (SSCNN) that utilizes full timing information in a 4D CNN, as described in [12].
@@
In the methods employing om2vec, we first pre-process the events using the trained 64-parameter om2vec model, to generate latent representations for the subsequent SSCNN and CNN networks. In Figure 6, we show the neutrino angular resolution (the angular difference between the true and predicted neutrino direction), as a function of the true neutrino energy. We find that the SSCNN models using the full timing information and the latent representations perform similarly, though as expected, the full 4D model is able to achieve a slightly better resolution. Notably, SSCNN sacrifices some performance to remain computationally feasible on higher-dimensional sparse data [26]. Thus, we observe that the dense CNN approach with latent representations performs just as well or slightly better than the SSCNN with full-timing information, depending on the true neutrino energy. This CNN approach is made possible by the fixed-length latent representations. Importantly, we note that this new flexibility allows for the easy adaptation of more powerful models, such as vision transformers [27], for angular reconstruction.
@@ 6 Conclusions and Future Directions
Availability: Source code, datasets, and pre-trained checkpoints for om2vec on Prometheus events is made available on GitHub. Additionally, as a future direction, om2vec is will be integrated into GraphNeT [28], an open-source deep learning pipeline widely utilized by various collaborations in the neutrino telescope community.